Can you please help me by showing me how to solve the following. It says solve the equation by completing the square. x^2 +8x=5 Thank you!
Answer by chessace(471)

(x+a)^2 = x^2 + 2a + a^2
So in this case a = 4
Complete the square by adding 4^2 to both sides
x^2 + 8x + 16 = 5 + 16
(x+4)^2 = 21
x+4 = +-sqrt(21)
x = -4 +-sqrt(21)